OpenStudy (anonymous):

Identify the property demonstrated: (5^3)^2 = 5^3 * 2 Power of a Power Property Quotient of Powers Property Product of Powers Property Power of a Product Property

OpenStudy (anonymous):

There is no dividing, so you can automatically cancel out B

OpenStudy (anonymous):

If it were being multiplied, it would be added That eliminates A

OpenStudy (anonymous):

Sorry I mean that means the answer is A XD Does that make sense?

OpenStudy (anonymous):

(a^3)(a^5) = a^(3+5) (a^3)^5 = a^(3)(5) = a^15